OpenStudy (anonymous):

Calculate the number of atoms in a 3.63 × 10^3 g sample of bromine.

OpenStudy (abb0t):

Always start with what you're given! \(\sf 3.63 \times 10^3 \times molar~mass~of~bromine\) which you can find on the periodic table. Look up, bromime, which is \(\sf \color{red}{Br}\) then, multiply by avogadro's number, which is: \(\sf \color{blue}{6.022 \times 10^{23} \frac{atoms}{mol}}\)

OpenStudy (anonymous):

Thanks!